At a recent rock concert, a dB meter registered 114 dB when placed 1.70 m in front of a loudspeaker on the stage.

(a) What was the power output of the speaker, assuming a hemispherical radiation of the sound and neglecting absorption in the air?
_____W 9.12W IS WRONG
(b) How far away would the intensity level be a somewhat reasonable 85 dB?
_____m
Answer #1

I (dB) = 10 109 (7) I. - 10-12 W/m2 114 = 10 169 (ton) I = 0.2511 W/m2 Now P-IA = I # 4*d? P = 4aId² P = 4*3.44* 0.2511W* 1.7